OpenStudy (anonymous):

Solve. Check for extraneous solutions. (x + 3)^1/2 - 1 = x sqrt 2x + 6 - sqrt x - 1 = 2 If you could show your work so I can follow along, that would be really helpful. Thank you!

OpenStudy (anonymous):

(x+3)^1/2=x+1 x+3=(x+1)^2 x+3=x^2+2x+1 x^2+2x+1-x-3=0 x^2+x-2=0 (x+2)(x-1)=0

OpenStudy (anonymous):

That makes SO much more sense now! Thank you so much!

OpenStudy (anonymous):

you are welcome